# Decomposing God Header File via Multi-View Graph Clustering

### v1

@@ IV Evaluation Set Up
    How well does our approach perform in terms of modularity?
@@ IV-A Datasets
Real-world God Header Files. To assess the practical performance of our approach, we also employed it on the typical God Header Files identified in section II-A (Table I). These files, chosen from manually confirmed God Header Files, possess the highest number of commits in their respective projects. They are characterized by an extensive amount of code elements, influencing hundreds of thousands of lines of code. The abundant commit histories of God Header Files is advantageous for us to estimate the reduction on recompilation after decomposition.
@@ IV-B Comparison Methods
In order to apply these methods to the task of God Header File decomposition, we reproduced their code similarity calculation module. In this process, we adhere to the original design of each comparison method, substituting all references to “attribute” and “method” with “code element”, and replace attribute access and method invocation with define-use relationships between code elements.
IV-C Metrics on Synthetic Data
To assess the accuracy of our approach and to answer RQ1, we measured each generated decomposition result based on its closeness to the expected result using a set of common metrics, which are described below.
@@ IV-C Metrics on Synthetic Data
    $MoJoFM(A,B)=(1-\frac{mno(A,B)}{\max(mno(\forall A,B))})\times 100\%,$ (7)
@@
We also utilize Accuracy (ACC) [34] and F1-score (F1) for evaluation, both calculated through cluster mapping provided by the Kuhn-Munkres algorithm [35].
IV-D Metrics on Real-world Data
To evaluate the performance on the real-world dataset and answer RQ2 and RQ3, we used the following metrics to assess the architecture and reduction on recompilation of the decomposition results.
@@ IV-D Metrics on Real-world Data
    $Q=\frac{1}{2m}\sum_{i,j}[A_{ij}-\frac{k_{i}k_{j}}{2m}]\delta(c_{i},c_{j}),$ (8)
@@ V-A RQ1: Accuracy
To assess the accuracy of the decomposing methods, we applied the three comparison methods along with our approach to the synthetic dataset. Since there is an expected number of subfiles for each synthetic God Header File to decompose, we adjusted the hyperparameters of each method to ensure that the decomposition results produce precisely that number of subfiles. We also conducted an ablation study on the synthetic dataset by removing the graph coarsening phase and the multi-view graph clustering phase (using K-means instead) respectively to assess their influence. The performance of each decomposing method and the results of our approach under each setting is presented in Table II.
@@
Figure 4: Modularity of different methods on real-world God Header Files. Black points represent the decomposition results with cyclic dependencies.
@@
Summary 1: In comparison to existing methods, our approach achieves 11.5% higher accuracy on all synthetic files and demonstrates enhanced consistency across various software projects. Both the graph coarsening component and the multi-view graph clustering component make a notable contribution to the performance of our approach.
V-B RQ2: Modularity
We also employed the three God Class refactoring methods and our approach to the six real-world God Header Files. In this situation, the number of subfiles required for decomposition is uncertain, so we empirically configured the number within a range from 4 to 10 for all six files to investigate the performance. For each decomposition result, we calculated Modularity and plotted the values as functions of the number of subfiles for different God Header Files, as illustrated in Figure 4.
Our approach consistently yields the highest Modularity across almost all configurations, with an average improvement of 9.1% compared to the other methods. This indicates that our approach generates sub-header files that adhere closely to the principles of high cohesion and low coupling, thereby facilitating architectural redesign. For the file SDL_dynapi_overrides.h, Wang’s approach achieves comparable Modularity with ours. This is because only one type of relationship dominates in this file: there are no dependencies and the number of co-usage edges is an order of magnitude less than semantic edges. In this scenario, the multi-view graph clustering module of our approach could not leverage its advantage. Meanwhile, the values of Modularity exhibit significant variations among different files. For instance, guc.h achieves a remarkable 40%, while Network.h struggles to surpass 3%. This discrepancy could be attributed to the inherent complexity and interwoven nature of the code elements in Network.h. Such God Header Files are possibly beyond the capabilities of automatic refactoring methods, emphasizing the necessity of involving experts in the architectural redesign.
In addition to the improvement on cohesion and coupling, which is evident from the Modularity values, our results also present acyclic dependencies among the generated sub-header files, thanks to the cyclic dependency fixing component. In our experiment, each refactoring method is applied to 6 files across 7 different settings, totaling 42 cases. Among these, 15 results of Wang’s approach and 28 of Akash’s exhibited cyclic dependencies, rendering them infeasible (indicated by black points in Figure 4). Bavota’s approach, in contrast, produced results without cycles due to its connected subgraph based clustering algorithm. However, these results often performed poorly on Modularity.
@@ V-B RQ2: Modularity
To illustrate the cyclic dependency fixing process, we provide an example in Figure 5. Figure 5 presents the clustering result for settings.h (#subfiles=7) after the multi-view graph clustering phase. In this instance, there are three cycles: 3-6-4-5, 3-6-4, and 3-6-5. Following our cyclic dependency fixing algorithm, we prioritize addressing the longest cycle, which is 3-6-4-5. The heuristic search process yields the optimal solution: moving the code element “rdpSettings” from subfile 4 to subfile 6. This movement involves only one code element and attains the highest gain of 61, eliminating dependencies from 4 to 5 and from 6 to 4. As a result, subfile 4 is no longer involved in any cycles. The intermediate result after this movement contains only one cycle: 3-6-5. In the next step, the best move is to relocate the code element “ALIGN64” from 3 to 5 based on the gain of each choice. Then all cycles are eradicated, and the final clustering result is presented in Figure 5.
Summary 2: Our decomposed header files exhibit superior architecture. They are characterized by an average improvement of 9.1% on Modularity compared to the state-of-the-art comparison method, along with simplified acyclic dependencies.
V-C RQ3: Reduction on recompilation
@@
Summary 3: Our approach could reduce recompilation of real-world God Header Files by 15% to 60%, and save at most 60 hours for one God Header File during its historical evolution.
@@ VI Discussion
In this section, we discuss the limitations and future work of our approach.
When practically applying our approach, the expected number of subfiles should be assigned by users. However, it’s not an easy task to determine the appropriate number. We suggest that users apply our approach on various settings and make trade-offs between the resulting Modularity and recompilation reduction. If multiple results meet the requirements, we recommend selecting the one with the minimum number for a simplified decomposition. We will investigate how to suggest the expected number automatically in the future.
Besides, our cycle dependency fixing algorithm is designed to guarantee that dependencies among the decomposed files are acyclic. However, this stage may disrupt the structure of clustering results in case the dependency relationships are complex. In future work, we will investigate how to integrate the acyclic constraint into the clustering process. Although there exists researches on acyclic partition[37, 38, 28], our specific task presents unique challenges, as our graph contains both directed and undirected edges.
Moreover, the current version of our approach does not assign names to the decomposed sub-header files. Each sub-header file should ideally be provided with a file name or a brief descriptive comment that reflects the concepts and functionalities of the file for developers to comprehend its contents quickly. We will further explore the process of generating appropriate file names and concise descriptive comments for these decomposed sub-header files.
VII Threats to Validity
The results of our experiments may suffer from several threats to validity. We discuss internal validity and external validity respectively.
@@ VIII Related Work
VIII-A Header File Optimization
There are plenty of researches aiming at reducing build time brought by header files. Some of them optimize the header files themselves, by removing false code dependencies[39], modifying unnecessary include directives[40, 41], or replacing include directives with forward declarations[42]. Others focused on optimizing the build processes through cache[43, 44], precompilation[45] or detection of redundant compilation[46].
Besides, McIntosh et al. [13] proposed an approach to identify header file hotspots, which undergo frequent modifications and trigger long-time rebuild processes. It helps developers identify and optimize header file hotspots, leading to reduced build times and increased productivity.
@@ VIII-B God Class Refactoring
God Class is one of the most concerned code smells, referring to those complex classes that centralize the intelligence of the system [47]. Most of the existing researches addressed the task in terms of cohesion and coupling, and applied techniques like static analysis, metric-based methods, or deep learning-based approaches. Bavota et al. [5, 6, 7] proposed a series of static analysis based approaches that extract the God class by calculating the cohesion between two methods in the class. Akash et al. [9] proposed a metric-based approach, applying Latent Dirichlet Allocation (LDA) [48] to represent the methods by a distribution of topics. With the development of deep learning techniques, the utilization of Graph Neural Network (GNN) model-based approaches has commenced to demonstrate its efficacy in addressing this particular task. A recent approach by Akash et.al [10] exploited the usage of graph auto-encoder to learn a vector representation for each method in the class. We learned from how these works extract code relationships and incorporated build dependencies that is crucial in our task. We also utilize a GNN based graph clustering algorithm for improved decomposition.
VIII-C Graph Clustering
@@
Graph clustering, also known as community detection, is a fundamental problem in network analysis and has been extensively studied across various domains. Traditional algorithms are characterized by their reliance on mathematical and structural properties of graphs, such as spectral clustering [49], modularity-based clustering like Louvain [50], hierarchical clustering [51], etc. However, these algorithms often lack scalability and struggle to handle large and complex graphs effectively. In recent years, many deep learning-based graph clustering methods have been proposed. Such methods typically learn low-dimensional representations of graph nodes through Graph Neural Networks (GNNs) and then use them for clustering [52, 26, 53].
@@
However, challenges persist in terms of the heterogeneity in real-world complex graph data, giving rise to multi-view graph clustering, which combines information from various sources or feature subsets to improve clustering quality, as opposed to methods that rely on a single graph representation. These approaches encompass a wide spectrum, including multiple kernel clustering (MKC) [54, 55], subspace clustering [56, 57], NMF-based (non-negative matrix factorization) multi-view clustering [58], ensemble-based multi-view clustering [59], etc. Notable works such as O2MAC [22], MvAGC [23], MCGC [24] and DuaLGR [25] have achieved excellent performance. They enhances learning performance by capitalizing the consistency and complementarity among different views and thus improves generalization compared to single-view clustering.
IX Conclusion
@@
In this paper, we make aware of the problem of God Header File and propose an approach to automatically decompose God Header Files in C projects. We evaluate our approach on both synthetic God Header Files and real-world God Header Files. The results reveal that, in comparison to existing methods, our approach attains 11.5% higher accuracy and exhibits more consistent performance across different projects. When applied to real-world God Header Files, our decomposition results get better modularity and 15% to 60% reduction on recompilation. In future work, we will explore how to recommend the optimal number of decomposed files, generate the appropriate file name, and refactor all the related files automatically, facilitating code comprehension and maintenance in C projects.

### v2

@@ IV Evaluation Setup
    How well does our approach perform in terms of refining architecture?
@@ IV-A Datasets
Real-world God Header Files. To assess the practical performance of our approach, we also employed it on the typical God Header Files identified in section II (Table I). These files, chosen from manually confirmed God Header Files, possess the highest number of commits in their respective projects. They are characterized by an extensive amount of code elements, influencing hundreds of thousands of lines of code. The abundant commit histories of God Header Files is advantageous for us to estimate the reduction in recompilation after decomposition.
@@ IV-B Comparison Methods
We chose Bavota’s [8] and Wang’s [26] approaches for comparison as they represent the state of the art. We did not include the well-known JDeodorant [27, 28] because both Bavota’s and Wang’s studies claimed superior performance compared to it. Additionally, Akash’s [11] approach was selected due to its use of Graph Neural Networks (GNNs), aligning with our own approach which also employs a GNN-based clustering algorithm.
In order to apply these methods to the task of God Header File decomposition, we reproduced their code similarity calculation module. In this process, we adhere to the original design of each comparison method, substituting all references to “attribute” and “method” with “code element”, and replace attribute access and method invocation with define-use relationships between code elements. Then we adopt the settings of clustering algorithms exactly as described in the comparison works.
IV-C Metrics
To assess the accuracy of our approach and answer RQ1, we measured each generated decomposition result based on its closeness to the expected result using a set of common metrics, which are described below.
@@ IV-C Metrics
    $\small MoJoFM(A,B)=(1-\frac{mno(A,B)}{\max(mno(\forall A,B))})\times 100\%$ (8)
@@
  • Accuracy (ACC) [32] measures the proportion of correctly predicted cluster labels by building a mapping between predicted cluster labels and ground truth cluster labels through the Kuhn-Munkres [33] algorithm.
  • F1-score (F1) is also a metric for evaluating the clustering performance, which combines precision and recall.
To evaluate the performance on the real-world dataset and answer RQ2 and RQ3, we used the following metrics to assess the architecture and reduction in recompilation of the decomposition results.
@@
    $Q=\frac{1}{2m}\sum_{i,j}[A_{ij}-\frac{k_{i}k_{j}}{2m}]\delta(c_{i},c_{j})$ (9)
@@ V-A RQ1: Accuracy
In RQ1, we applied the three comparison methods along with our approach to the synthetic dataset. Since there is an expected number of subfiles for each synthetic God Header File to decompose, we adjusted the hyperparameters of each method to ensure that the decomposition results produce precisely that number of subfiles. The performance of each decomposing method and the results of our approach under each setting is presented in Table II.
@@
Summary for RQ1 In comparison to existing methods, our approach achieves 11.5% higher accuracy on all synthetic files and demonstrates enhanced consistency across various software projects. Both the graph coarsening component and the multi-view graph clustering component make a notable contribution to the performance of our approach.
V-B RQ2: Architectural benefits
We also employed the three God Class refactoring methods and our approach to the six real-world God Header Files to investigate the performance. In this situation, the number of subfiles required for decomposition is uncertain, so we empirically configured the number within a range from 4 to 10 for all six files , as developers won’t prefer a larger number in practice. In this section, we will discuss the architectural benefits of our approach from the perspective of acyclic dependencies and higher modularity among the sub-header files of the decomposition results.
Figure 4: Modularity of different methods on real-world God Header Files. Black cross markers represent the decomposition results with cyclic dependencies.
In Figure 4, we calculated modularity for each decomposition result and plotted the values as functions of the number of subfiles for different God Header Files. The decomposition results that suffered cyclic dependencies, which may lead to compilation errors, are marked by black cross markers. In our experiment, each refactoring method is applied to 6 files across 7 different settings, totaling 42 cases. Among these, 15 results of Wang’s approach and 28 of Akash’s exhibited cyclic dependencies, rendering them infeasible. Bavota’s approach, in contrast, produced results without cycles due to its connected subgraph based clustering algorithm. However, these results often performed poorly on modularity. As for our approach, the cyclic dependency fixing module ensures acyclic dependencies among the decomposed sub-header files, and the multi-view graph clustering module guarantees high-level modularity, resulting in a simpler and clearer architecture.
To be specific, our approach consistently yields the highest modularity across almost all configurations, with an average improvement of 9.1% compared to the best comparison method. This indicates that our approach generates sub-header files that adhere closely to the principles of high cohesion and low coupling, thereby facilitating architectural redesign. For the file SDL_dynapi_overrides.h, Wang’s approach achieves comparable modularity with ours. This is because only one type of relationship dominates in this file: there are no dependencies and the number of co-usage edges is an order of magnitude less than semantic edges. In this scenario, the multi-view graph clustering module of our approach could not leverage its advantage. Meanwhile, the values of modularity exhibit significant variations among different files. For instance, guc.h achieves a remarkable 40%, while Network.h struggles to surpass 3%. This discrepancy could be attributed to the inherent complexity and interwoven nature of the code elements in Network.h. Such God Header Files are possibly beyond the capabilities of automatic refactoring methods, emphasizing the necessity of involving experts in the architectural redesign.
@@ V-B RQ2: Architectural benefits
To illustrate the cyclic dependency fixing process, we provide an example in Figure 5. Figure 5(a) presents the clustering result for settings.h (#subfiles=7) after the multi-view graph clustering phase. In this instance, there are three cycles: 3-6-4-5, 3-6-4, and 3-6-5. Following our cyclic dependency fixing algorithm, we prioritize addressing the longest cycle, which is 3-6-4-5. The heuristic search process yields the optimal solution: moving the code element “rdpSettings” from subfile 4 to subfile 6. This movement involves only one code element and attains the highest gain of 61, eliminating dependencies from 4 to 5 and from 6 to 4. As a result, subfile 4 is no longer involved in any cycles. The intermediate result after this movement contains only one cycle: 3-6-5. In the next step, the best move is to relocate the code element “ALIGN64” from 3 to 5 based on the gain of each choice. Then all cycles are eradicated, and the final clustering result is presented in Figure 5(b).
Summary for RQ2 Our decomposed sub-header files exhibit superior architecture. They are characterized by an average improvement of 9.1% on modularity compared to the state-of-the-art comparison method, along with simplified acyclic dependencies.
V-C RQ3: Reduction in recompilation
@@
Summary for RQ3 Decomposing God Header Files into subfiles using our approach can significantly reduce recompilation cost, ranging from 15% to 60%. Our method generally achieves the lowest recompilation cost compared to other approaches and can save up to 60 hours for a single God Header File over its historical evolution.
@@ VI Discussion
In this section, we discuss the application, limitations and threats to validity of our approach.
VI-A Application and Limitation
Our approach aims to provide an improved decomposition solution for God Header Files. In the experiments, the appropriate number of sub-header files can be suggested by balancing modularity and recompilation reduction. However, users may feel the recommended number of subfiles a bit excessive in application, as they may prefer a more concise file structure. Therefore, we allow users to assign the number of subfiles when practically applying our approach in their projects. We will also investigate how to suggest the best number automatically in the future.
In our approach, a cycle dependency fixing algorithm is designed to guarantee that dependencies among the decomposed files are acyclic. However, this stage may disrupt the structure of clustering results in case the dependency relationships are complex. In future work, we will investigate how to integrate the acyclic constraint into the clustering process. Although there exists researches on acyclic partition [35, 36, 24], our specific task presents unique challenges, as our graph contains both directed and undirected edges.
The current version of our approach does not try to assign a good name for the decomposed sub-header file. Ideally, each sub-header file should be provided with a file name or a brief descriptive comment that reflects the concepts and functionalities of the file for developers to comprehend its contents quickly. We will further explore the process of generating appropriate file names and concise descriptive comments for these decomposed sub-header files.
VI-B Threats to Validity
Our experiment results may suffer from several threats to validity. We discuss internal validity and external validity respectively.
@@ VII Related Work
VII-A Header File Optimization
There are plenty of researches aiming at reducing build time brought by header files. Some of them optimize the header files themselves, by removing false code dependencies [37], modifying unnecessary include directives [38, 39], or replacing include directives with forward declarations [40]. Others focused on optimizing the build processes through cache [41, 42], precompilation [43] or detection of redundant compilation [44]. Besides, McIntosh et al. [12] proposed an approach to identify header file hotspots, which undergo frequent modifications and trigger long-time rebuild processes. It helps developers identify and optimize header file hotspots, leading to reduced build times and increased productivity.
@@ VII-B God Class Refactoring
God Class is one of the most concerned code smells, referring to those complex classes that centralize the intelligence of the system [45]. While some researchers question the necessity of refactoring God Classes [46, 47], numerous studies have tackled this issue through semi-automatic [48] or automatic approaches, primarily focusing on cohesion and coupling analysis. These studies utilize techniques such as static analysis, metric-based methods, and even deep learning approaches. Bavota et al. [6, 7, 8] proposed a series of static analysis based approaches that extract the God class by calculating the cohesion between two methods in the class. Akash et al. [10] proposed a metric-based approach, applying Latent Dirichlet Allocation (LDA) [49] to represent the methods by a distribution of topics. With the development of deep learning techniques, the utilization of Graph Neural Network model-based approaches has commenced to demonstrate its efficacy in addressing this particular task. A recent approach by Akash et.al [11] exploited the usage of graph auto-encoder to learn a vector representation for each method in the class.
We learned from how these works extract code relationships and incorporated build dependencies that is crucial in our task. We also utilize a GNN based graph clustering algorithm for improved decomposition.
VII-C Graph Clustering
@@
Graph clustering, also known as community detection, is a fundamental problem in network analysis and has been extensively studied across various domains. Traditional algorithms are characterized by their reliance on mathematical and structural properties of graphs, such as spectral clustering [50], modularity-based clustering like Louvain [51], hierarchical clustering [52], etc. However, these algorithms often lack scalability and struggle to handle large and complex graphs effectively. In recent years, many deep learning-based graph clustering methods have been proposed. Such methods typically learn low-dimensional representations of graph nodes through Graph Neural Networks and then use them for clustering. Some notable methods, such as GraphSAGE [53], Graph Convolutional Networks (GCNs) [22], and Variational Graph Autoencoders (VAGE) [54], etc., outperform traditional clustering methods.
@@
Despite these advancements, challenges remain due to the heterogeneity of real-world complex graph data, leading to the development of multi-view graph clustering. These works combine information from various sources or feature subsets. By leveraging the consistency and complementary of different views, they exhibit superior effectiveness and generalization compared to single-view clustering. These approaches encompass a wide spectrum, including multiple kernel clustering (MKC) [55, 56], subspace clustering [57, 58], NMF-based (non-negative matrix factorization) multi-view clustering [59], ensemble-based multi-view clustering [60], etc. Notable works such as O2MAC [61], MvAGC [62], MCGC [63] and DuaLGR [21] have achieved excellent performance. In our work, we employ DuaLGR due to its effectiveness with non-homophilous edges, dynamic weight assignment to address inconsistencies across projects, and scalability to incorporate additional relationships.
VIII Conclusion
@@
In this paper, we make aware of the problem of God Header File and propose an approach to automatically decompose God Header Files in C projects. We evaluate our approach on both synthetic God Header Files and real-world God Header Files. The results reveal that, in comparison to existing methods, our approach attains 11.5% higher accuracy and demonstrates more consistent performance across different projects. When applied to real-world God Header Files, our decomposition results exhibit better architecture, evidenced by higher modularity and acyclic dependencies, and achieve a 15% to 60% reduction in recompilation. In future work, we will explore how to recommend the optimal number of decomposed files, generate the appropriate file name, and refactor all the related files automatically, facilitating code comprehension and maintenance in C projects.
Acknowledgment
This work is supported by the National Key Research and Development Program of China under Grant No.2023YFB4503803.